Problem 4-22 Sustainable Growth Rate [LO3] Cambria, Inc., had equity of $200,000 at the beginning of the year. At the end of the year, the company had total assets of $355,000. During the year the company sold no new equity. Net income for the year was $42,000 and dividends were $6,000. What is the sustainable growth rate for the company? (Do not round intermediate calculations and round your final answer to 2 decimal places. (e.g., 32.16) Sustainable growth rate What is the sustainable growth rate if you use the formula ROE x b and beginning of period equity? (Do not round intermediate calculations and round your final answer to 2 decimal places. (e.g., 32.16)) Beginning of period equity What is the sustainable growth rate if you use end of period equity in this formula? (Do not round intermediate calculations and round your final answer to 2 decimal places. (e.g., 32.16)) End of period equity
Answer #1

Beginning Equity = $200,000

Ending Equity = Beginning Equity + Net Income - Dividends
Ending Equity = $200,000 + $42,000 - $6,000
Ending Equity = $236,000

Dividend Payout Ratio = Dividends / Net Income
Dividend Payout Ratio = $6,000 / $42,000
Dividend Payout Ratio = 0.1429

Retention Ratio, b = 1 - Dividend Payout Ratio
Retention Ratio, b = 1 - 0.1429
Retention Ratio, b = 0.8571

Answer a.

Return on Equity, ROE = Net Income / Ending Equity
Return on Equity, ROE = $42,000 / $236,000
Return on Equity, ROE = 0.1780 or 17.80%

Sustainable Growth Rate = [ROE * b] / [1 - ROE * b]
Sustainable Growth Rate = [0.1780 * 0.8571] / [1 - 0.1780 * 0.8571]
Sustainable Growth Rate = 0.15256 / 0.84744
Sustainable Growth Rate = 0.1800 or 18.00%

Answer b.

Return on Equity, ROE = Net Income / Beginning Equity
Return on Equity, ROE = $42,000 / $200,000
Return on Equity, ROE = 0.2100 or 21.00%

Sustainable Growth Rate = ROE * b
Sustainable Growth Rate = 0.2100 * 0.8571
Sustainable Growth Rate = 0.1800 or 18.00%

Answer c.

Return on Equity, ROE = Net Income / Ending Equity
Return on Equity, ROE = $42,000 / $236,000
Return on Equity, ROE = 0.1780 or 17.80%

Sustainable Growth Rate = ROE * b
Sustainable Growth Rate = 0.1780 * 0.8571
Sustainable Growth Rate = 0.1526 or 15.26%